4. (a) State the Bio-Savart law. (1 mark) (b) Use the Bio-Savart law to calculate the magnetic field at a point, P, a distanc question b
Answer #1

b) see the diagram:

dx > Khaota -L D AL dan figol

Refer to the fig. 1:

Considering an element at a distance x from center of length dx.

sin\theta = \frac{L}{\sqrt{r^{2}+L^{2}}}

x=r\tan\alpha

\Rightarrow dx=r\sec^{2}\alpha*d\alpha------(i)

also, coa\alpha = \frac{r}{\sqrt{x^{2}+r^{2}}}

Apply Biot-Savart's Law:

magnetic field due to the element is

dB = \frac{\mu_{o}i*dx*sin(90^{\circ}-\alpha)}{4\pi (\sqrt{r^{2}+x^{2}})^{2}}

The net magnetic field is

B = \int_{-\theta}^{\theta}\frac{\mu_{o}i*dx*cos(\alpha)}{4\pi (\sqrt{r^{2}+x^{2}})^{2}}

= \int_{-\theta}^{\theta}\frac{\mu_{o}i*cos(\alpha)*r*\sec^{2}\alpha*d\alpha}{4\pi (\sqrt{r^{2}+x^{2}})^{2}}

= \int_{-\theta}^{\theta}\frac{\mu_{o}i*cos(\alpha)*r^{2}*\sec^{2}\alpha*d\alpha}{4\pi r(\sqrt{r^{2}+x^{2}})^{2}}

= \int_{-\theta}^{\theta}\frac{\mu_{o}i*cos(\alpha)*cos^{2}\alpha*\sec^{2}\alpha*d\alpha}{4\pi r}

= \frac{\mu_{o}i}{4\pi r}\int_{-\theta}^{\theta}cos(\alpha)*d\alpha

= \frac{\mu_{o}i}{4\pi r}*2sin\theta

= \frac{\mu_{o}i}{2\pi r}*\frac{L}{\sqrt{r^{2}+L^{2}}} [answer]

Refer to the fig. 2:

1596230595429_WhatsAppImage2020-08-01at2

The length of each side is 2L.

the distance of the center from the side of the square is r = L.

The magnetic field due to one side is

B=\frac{\mu_{o}i}{2\pi r}*\frac{L}{\sqrt{r^{2}+L^{2}}} [using the result above]

As the direction of the magnetic field due to all the sides at the center is the same and has the same magnitude.

Therefore, the net magnetic field at the center is B_{net}=4B=4*\frac{\mu_{o}i}{2\pi *L}*\frac{L}{\sqrt{L^{2}+L^{2}}}

  =1*\frac{\mu_{o}i}{\pi}*\frac{1}{L\sqrt{2}}

  =\frac{\mu_{o}i}{\sqrt{2}\pi L} [answer]